

********************************************************************
UARS WEEKLY STATUS REPORT - 25 Sep 1998
********************************************************************
***********
SOLSTICE
***********
With the exception of the ~8 hour unscheduled off period due to the spacecraft
under-voltage condition on Thursday (98267), SOLSTICE continues to operate
normally, collecting solar data on all available orbital tracking periods and
stellar calibration observations, whenever possible.
***********
SUSIM
***********
Nothing received.
***********
PEM
***********
The PEM zenith boom instruments were active from September 15 through
September 19, when the zenith instruments were deactivated and the PEM/AXIS
was activated. The PEM/AXIS was deactivated on September 23 due to low
spacecraft power. During these activation periods, PEM produced both electron
and ion energy deposition from in situ measurements when the zenith
spectrometers were active and produced energy deposition from electrons,
inverted from measured X-rays obtained in 16 pixels, from the Atmospheric
X-ray Imaging Spectrometer.
GSFC has proposed an OBC operation plan for the activation of PEM boom
instruments during the daylight portions of every orbit. This OBC operation
plan is accepted by PEM. PEM has proposed a series of test procedures for
activation and is waiting response from GSFC. PEM is also waiting for GSFC
scheduling and implementation of its operation test procedures.
***********
HALOE
***********
HALOE continues successful operations with all engineering parameters being
nominal and the science data continues to be of excellent quality. HALOE
was off for approximately 5.5 hours due to the low voltage situation on
September 24th. HALOE is scheduled to be turned off near the end-of-day
September 26, 1998 for the upcoming yaw.
***********
MLS
***********
MLS is scheduled to begin operations on Oct 6 and operate though Oct 10. The
MLS team will operate the instrument in a mode for diagnosing the reoccurance
of slips in the antenna scan mechanism, and to attempt to find a method of
operating the instrument which minimizes the impact of these slips on
scientific data.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
Routine processing of instrument Daily Activity Plans continues
with no other problems. The MPG continues to assist the Project
Scientist with the scheduling of instrument operations.
***********
CDHF
***********
The CDHF has received a total of 420 CD-ROM requests as of 25-September-1998.
There were seven orders received and shipped.
The CDHF processed current day data through level 3 for all active
instruments. Additionally, CDHF personnel reprocessed: (R)= Requested
(C) = Complete, (F)= Failed (L) = Days remaining to reprocess.
52 days HAL (4.80) PROD (L1) (1873R-1500C-36F=337L)
31 days HAL (4.80) PROD (L3) (1873R-858C=1015L)
87 days SOL (4.62) PROD (L2-L3) (2349R-1601C=748L)
8 days SUS (7.00) ENG (65R-65C=0L)
2 days WIN (5.10) ENG (4R-4C=0L)
The CDHF loaded July DAAC data into the DAAC Transfer Tables.
The CDHF executed and monitored a SOLSTICE test day during standalone time.
The test permited investigation of a wallclock time inflation problem
reported by the SOLSTICE Representative.
CDHF pending software upgrades: IDL 5.1, OpenIngres.
***********
FOT
***********
REPORTING PERIOD
This report covers 12 September 1998 (Orbit 38265, GMT Day 255) through 18
September 1998 (Orbit 38369, GMT Day 261).
SPACECRAFT OPERATION
The observatory is now in Reverse Flight (Northern Hemisphere viewing) and
is performing nominally under two-battery operations. The instrument
operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
MLS 257/2252 38309 Off
ACRIM 257/2252 38309 On
SUSIM 258/0010 38310 On
PEM 258/0512 38313 Zeps On
WINDII 259/0014 38325 On
The beta angle ranged between 52.4 degrees to 41.5 degrees for this report
period. The beta angle is now decreasing toward a minimum angle of -2.0
degrees for the yaw on 29 September 1998 (DOY 272).
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 23.8 V end-of-night (EON) load bus voltage.
Battery temperatures are stable with a temperature delta between Battery 2
and Battery 3 of 0.9 to 1.2 degrees C. Additional battery performance data
for each battery on 11 September 1998 and 18 September 1998 are provided in
Appendix D.
The clock error ranged from 11.2 to -11.3 msec during this report period.
Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT for
this report. A summary is listed below and the details of ALERT and Out-
Of-Limits occurrences are listed in Appendix B.
% ALERTS = 00
% OUT-OF-LIMITS = 03
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
% FOT 04
% MLS 01
% ACRIM 01
% PEM 01
% SUSIM 01
% WINDII 01
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
% FOT 01
One (1) Anomaly Report remained OPEN at the end of this report period (see
Appendix E).
TRANSPONDER FREQUENCIES:
Transponder A: 98/210 - 2287.496941
Transponder B: 98/211 - 2287.499161
GROUND SYSTEM OPERATION
No operational support problems occurred this period. Anomaly Reports are
included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from
mission start through UTC Day 98/260 show the observatory data loss to be
21 hours, 06 minutes, 41.140 seconds (an increase of 1 minute and 39.136
seconds since last report period). This is a 0.0345 percent data loss
which equals a 99.9655 percent data capture for the mission.
The increase of 1 minute and 39.136 seconds occurred over three days:
7 1:27.776 minutes on Day 254 due to line errors and data gap,
7 10.336 seconds on Day 255 due to data gap, and
7 1.024 seconds on Day 257 insufficient overlap and hit in the data.
The cause of the data gap(s) is under investigation to exclude tape
recorder involvement.
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
FOT continues to improve a power profile for UARS based on the current
status of the Modular Power System.
The possible impacts of the Leonid meteorite shower in November 1998 are
being investigated.
YEAR 2000 OPERABILITY SUPPORT
UARS Test and Training Simulator (UTTS) - No change.
User Planning System (UPS) - No change.
Multi Satellite Operations Control Center (MSOCC) - The FOT attended a
meeting on 17 September to discuss the Year 2K compatibility of the POCC AP
software, as well at the UARS Test and Training Simulator (UTTS). The Year
2K POCC/AP test team had several questions related to FOT. In addition,
the test team pointed out a few Year 2K related AP software problems, and
asked the FOT whether or not they could live with the problems. The FOT
agreed to investigate the test teamUs findings and respond by early next
week.
Command Management System (CMS) - The FOT successfully processed a Year 2K
User Antenna View (UAV) file on the back-up CMS this week. The file
appeared to process without any problems, and a report of the generated ODB
depicted a proper roll-over from 1999 to 2000. The FOT also received a
Year 2K PSAT file, but cannot process it until the corresponding S/C
Orbital Events file is received (due to a CMS processing dependency).
Finally, the FOT received 3 corrected UARS OBC ephemeris files. However,
the new files still did not contain a full 72 hours worth of data and hence
could not be processed by the CMS. FDF was notified of the problem (once
again) and stated that they would correct these files and re-send them next
week.
Flight Dynamics Facility (FDF) - The FOT is still waiting on a big portion
of the planning aid data files. The pending files are either still being
generated by FDF or are hung up at the test FORMATS system in Greentech 4.
As soon as the FOT gets a complete set of planning aids, Year 2K testing
can proceed. However, it should be noted that due to the lack of FDF
supplied planning aids, the CMS test team will be under a severe time
constraint to meet the September 30th deadline.
OBC PMON FLIGHT SOFTWARE UPDATE
Flight Software - Complete.
Ground Support Software - Complete.
Verification Testing - Complete.
Uplink - Supporting data (tapes, etc.) for the UARS Test and Training
Simulator (UTTS) for scenario generation were accumulated on 14 and 15
September. The scenario, executed on 16 September, provided a more
accurate representation of what the telemetry should look like before,
during and after the software uplink. The simulation of the uplink was
successful. This simulation will be repeated a day prior to uplink.
IP NETWORK TESTING
Space Network (SN) - No change.
Testing with the NCC 98 system to support routine flight operations is
being planned.
Deep Space Network (DSN) and Ground Spaceflight Tracking and Data Network
(GSTDN) - Performed a DSN data flow from Johnson Space Center (JSC) using
pre recorded UARS data. 1kbps data and 32kbps I channel real-time data
were received. Command was not possible because the AP can not generate
the correct destination code for the JSC simulator - not in UARS AP
software. JSC attempted to send 32k OBC dump data but was unsuccessful.
JSC may not be able to simulate that type of data.
UARS PROPULSION MODULE B/D THRUSTER LINE FILL
The B/D thruster refill effort continues. The Eaton Latch Valve test
report is still being reviewed by the UARS FOT.
UARS RE-ENGINEERING - TELEMETRY PROCESSING
Flight operations is providing support to the re-engineering effort to
process the UARS telemetry on the TOMS/EP control center computer system.
Two MSOCC terminals were removed from the operations console on 10
September and replaced with two multi-purpose terminals driven by the re-
engineering system. The re-engineering system is being RrunS in parallel
to the MSOCC system for evaluation purposes.
UARS COMMAND RE-HOST PHASE 1 DEMONSTRATION
Flight operations continues to support the update of the Lockheed Martin
SCS-21 control center system. Development continues on the system. A
Phase 1 feasibility demonstration of the SCS-21 system supporting UARS is
scheduled for 22 September 1998.
MEETINGS
The FOT attended a meeting on 17 September 1998 to discuss the Year 2K
compatibility of the POCC AP software, as well at the UARS Test and
Training Simulator (UTTS).
UARS and TOMS/EP participated in the GSFC On-Orbit Operations Independent
Study Team (IST) interviews with the GSFC operational mission teams on 17
September 1998. Presentations in handout form were prepared according to
the study teamUs outline.
The next battery telecon is scheduled for 22 September 1998 at 1:00 pm.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw Around Maneuver is scheduled for scheduled for 29
September 1998.
The United States Naval Observatory (USNO) issued a Time Service
announcement advising that the International Earth Rotation Service (IERS)
will retard UTC time by 1.0 second at the end of December 1998. This means
the addition of one leap second as the last second of 1998.
The next annular solar eclipse will occur on 16 February 1999 over the
Indian Ocean, Australia and the South Pacific Ocean.
PERSONNEL CHANGES
Kathy Palmer, Flight Operations Director, submitted her resignation
effective 24 September 1998.
Deryck Chetram, Flight Operations Director, submitted his resignation
effective 25 September 1998.
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 18 September 1998
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ -------
38304 257/1529 98-342 OBC Clk Adj fr 20.6 to 0.00 msec/day
38309 257/2252 98-343 MLS Power OFF
38309 257/2252 98-344 ACRIM Power ON
38310 258/0010 98-345 SUSIM Power ON
38313 258/0512 98-346 PEM ZEPS Power On
38325 259/0014 98-347 WINDII Power ON
38330 259/0751 98-348 OBC Clk Adj fr 0.00 to 20.6 msec/day
38333 259/1306 98-349 RTS SUSIM Heater Cycle
38367 261/1919 98-350 OBC Clk Adj fr 20.6 to 0.00 msec/day
_______________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 18 September 1998
MNEMONIC STATE DESCRIPTION ORBITS
--------- ----- ------------ ------
QTSISAIF Y-HI SSPP Baseplate by SISA I/F 38280-317
QTSSPPMX Y-HI SSPP -X Thermostat 38282
NBRECBI Y-HI NBTR B Current 38282,97,303,330
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
--------- ------------ ------
None this report period
_______________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 18 September 1998
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ---- -----------------
38304 257/15:32 (14 Sept 98) from 20.6 to 00.0 msec/day
38330 259/07:59 (16 Sept 98) from 00.0 to 20.6 msec/day
38367 261/19:28 (18 Sept 98) from 20.6 to 00.0 msec/day
_______________________________________________________________
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 18 September 1998
11 Sept 1998 - GMT Day 254 Beta = 52.4 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 +728.0/+728.0 -3.05/-4.31 +0.0/ +0.0 18.1 0.000 0.0
2 +39.2/-11.2 +5.79/+4.51 +26.0/-16.0 25.4 1.024 21.8
3 +44.8/-11.2 +4.51/+3.23 +25.2/-16.0 25.3 1.027 21.3
18 Sept 1998 - GMT Day 261 Beta = 41.5 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 +728.0/+728.0 -3.30/-2.73 +0.0/ +0.0 18.1 0.000 0.0
2 +67.2/ 0.0 +5.71/+6.44 +28.8/-16.0 24.6 1.022 22.5
3 +33.6/-22.4 +4.77/+5.47 +28.4/-16.8 24.6 1.023 22.1
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- -------- ----------
no changes (V/T 5)
APPENDIX E
UARS ANOMALY REPORTS
Week ending 18 September 1998
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
98-060 FOT 38313 258/0512 N/A TAC Crashed
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
98-059 FOT 38224 253/14:10 20186 NO CMD/IP SUPPORT
98-060 FOT 38313 258/05:12 TAC 8 CRASH
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------- ------ ----- ---- --- -----------------
None this report period
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
98-057 FOT Sept 98 See list 14170 GENERIC LATE ACQ